A student makes bracelets and neckplaces starting on August 1st. She finishes a necklace every 6 days and a bracelet every 8 days. What is the first date on which the student will finish both a necklace and a bracelet?

Found 3 solutions by ikleyn, Theo, MathTherapy:
Answer by ikleyn(52776)

.

First date on which the student will finish both a necklace and a bracelet in the same day is August 24.

24 is the Least Common Multiple of the numbers 6 and 8.


Solved.


//////////////


Ignore the answer  ( " August, 25 " ) by  @Theo,  since it is not consistent with the problem.


=================


See the lesson
    - Light flashes on a Christmas tree and a Least Common Multiple
in this site and find there other similar solved problems.

Reading this lesson will help you to learn a conception of a Least Common Multiple in full.



Answer by Theo(13342)

if they start august 1st, the first date that they will finish both on the same date would be 24 days after the first = august 25.
for the necklase, august 1 + 6 = 7 + 6 = 13 + 6 = 19 + 6 = august 25.
for the bracelet, august 1 + 8 = 9 + 8 = 17 + 8 = august 25.

how did i get 24?
multiples of 6 are 6,12,18,24,.....
multiples of 8 are 8,16,24,.....
lowest common multiple is 24.


Answer by MathTherapy(10551)

A student makes bracelets and neckplaces starting on August 1st. She finishes a necklace every 6 days and a bracelet every 8 days. What is the first date on which the student will finish both a necklace and a bracelet? 
Correct answer:  

On the 24th day of month, the faster student will have finished 4 necklaces, while the slower will have finished 3 bracelets.